### v2

@@ 4 Improved Approximations for Two Agents
In light of our results in the previous sections, which show that the surplus guarantees achievable in several well-studied multi-parameter auction settings with many agents are no worse than the surplus guarantees achievable in the canonical single-item setting with many agents, a natural question is whether there exists settings which are more challenging from an approximation standpoint than the single-item setting. To further explore this question, we now turn to the interesting special case of two agents. In previous work of [hartline2008optimal], it was demonstrated that there exists a mechanism for two agents and a single item which achieves surplus which is a $3/2$-approximation to the optimal social welfare [^5] More precisely, they demonstrate that their mechanism achieves a $3/2$-approximation to a weaker benchmark $\mathcal{G}$ which we discuss in Subsection 4.1. The proof of this guarantee, however, implies a $3/2$-approximation to the stronger benchmark of optimal social welfare, which is of interest to the majority of our paper.. We generalize this result with Mechanism 3 which, as we demonstrate in Theorem 16, obtains a $3/2$-approximation guarantee in the case of two agents with arbitrary monotone valuation functions over an items set $\mathcal{I}$.
Input: An items set $\mathcal{I}$, a valuation profile of two agents $v_{1},v_{2}$
@@ 2 With probability $1/3$: Allocate the grand bundle of items to bidder $1$ and charge nothing
Consider two agents drawing their values i.i.d. from a standard exponential distribution. As demonstrated by [hartline2008optimal], the mechanism which maximizes the surplus for such a setting randomly allocates the item to one of the two agents and the expected surplus of such a mechanism is $1$. On the other hand, the expected social welfare is equal to the expectation of the maximum of two random variables $X_{1}$ and $X_{2}$ drawn independently from an exponential distribution. We may compute this as
@@
4.1 Alternative Benchmarks
Recall that the mechanisms in Sections 2 and 3 achieve surplus guarantees which are optimal approximations to the social welfare. However, these guarantees are asymptotic and logarithmic, raising the question of whether or not there exist tighter benchmarks which can better separate the performance of two mechanisms. As all prior-free mechanisms are anonymous (i.e., cannot treat differently agents which are, a priori, identical), we would like to compare the performance of a prior-free mechanism to the surplus achieved by the best anonymous, truthful mechanism. This observation led to the development of the benchmark $\mathcal{G}(\vec{v})$ in [hartline2008optimal]. Formally, $\mathcal{G}(\vec{v})=\sup_{F}\text{Opt}_{F}(\vec{v})$ where $\text{Opt}_{F}(\vec{v})$ is the surplus of the Bayesian optimal mechanism when all bidders draw their value from a known distribution $F$ and the realization of values is $\vec{v}$. hartline2008optimal provide a prior-free mechanism which obtains a $O(1)$-approximation to $\mathcal{G}$ in the case of unit-demand bidders and identical items. In addition, they raise an open question of determining the best-possible approximation ratio when the benchmark is $\mathcal{G}$ “even in the two agent, one unit special case” [hartline2008optimal]. We resolve this open question below.
@@ 4     With the remaining probability allocate the item to bidder $2$ and charge bidder $2$ a price of 00
In Theorem 18 we demonstrate the approximation ratio of Mechanism 4 and in Theorem 19 we prove that the mechanism is truthful in expectation and ex-post individual rational.
@@
Mechanism 4 obtains surplus utility that is a $5/4$-approximation to $\mathcal{G}(\vec{v})$
@@
Index the bidders in decreasing order of value such that we have $v_{1}\geq v_{2}$. As demonstrated in [hartline2008optimal], in the special case of two agents and a single item, $\mathcal{G}(\vec{v})$ is equal to the better of performing a Vickrey auction or random allocation. In other words, we have that $\mathcal{G}(\vec{v})=\max\left\{v_{1}-v_{2},\frac{v_{1}+v_{2}}{2}\right\}$ [^6] We note that in the proof of Proposition 4.1 in [hartline2008optimal], after arguing that $\mathcal{G}(\vec{v})$ is the better of a Vickrey auction or random allocation in the case of two agents, they incorrectly compute that $\mathcal{G}(\vec{v})=\max\left\{v_{1}-\frac{v_{2}}{2},\frac{v_{1}+v_{2}}{2}\right\}$. This leads them to give an incorrect lower bound of $4/3>5/4$ on the approximation achievable by any mechanism.
@@
Theorem 19.
Mechanism 4 is truthful-in-expectation and ex-post individually rational.
Proof.
Recall that in single-parameter environments, a mechanism is truthful if it has a monotone allocation rule and charges payments corresponding to the identity given in [myerson1981optimal]. By observation, it is clear that Mechanism 4 is monotone, i.e., the probability a bidder is allocated a good is (weakly) increasing with her bid (holding the other bid constant). As such, it remains to verify that the prices charged by the mechanism are the threshold payments (i.e., those implied by the payment identity of myerson1981optimal). Consider varying the value $v_{i}$ of bidder $i$ when the value of the other bidder $j$ is fixed to $v_{j}$. When $0\leq v_{i}<\frac{1}{3}v_{j}$ the probability that bidder $i$ receives the good is $0.2$ and she pays 00, as expected by the payment identity. When $\frac{1}{3}v_{j}\leq v_{i}\leq 3v_{j}$ the probability that $i$ receives the good is $0.5$. As such, by the payment identity, $i$ should be charged $\frac{1}{3}v_{j}\cdot(0.5-0.2)=\frac{1}{10}v_{j}$ and this is precisely the price the mechanism charges (normalized by the probability to be allocated which is half). Finally, when $3v_{j}<v_{i}$ the probability that $i$ receives the good is $0.8$. But then, by the payment identity, $i$ should be charged $3v_{j}\cdot(0.8-0.5)+\frac{1}{10}v_{j}=v_{j}$ which is precisely what the mechanism charges (again, normalized by the allocation probability $\frac{4}{5}$). ∎
@@
Our proof proceeds very similarly to the proof of Proposition 4.1 in [hartline2008optimal] (essentially, correcting an arithmetic error which propagates therein). Consider two agents drawing their values i.i.d. from a standard exponential distribution. Again, the mechanism which maximizes the expected surplus for such a setting randomly allocates the item to one of the two agents and the expected surplus of such a mechanism is $1$. On the other hand, the expected value of $\mathcal{G}(\vec{v})$ is
@@
$=\frac{5}{4}.\qed$
Discussion and Future Directions
In this paper, we study mechanism design for maximizing consumer surplus. Our work is the first to achieve a sub-linear approximation to the welfare in multi-parameter settings, and we obtain asymptotically optimal guarantees for various settings including $O(\log(n))$ approximation for unit-demand and multi-unit submodular agents, $O(1+\log\frac{n}{m})$ for Bayesian unit-demand setting, and $O(\log(n))$ for agents with concave valuations over divisible goods. We also resolve open questions from [hartline2008optimal].
Beyond Unit-Demand and Multi-Unit.
We next discuss how to adapt Mechanism 1 for a broader class of valuations. Consider the case where agents’ valuations are weighted matroid valuation functions under known matroids. In other words, each agent $i\in\mathcal{N}$ is associated with a (publicly known to the auctioneer) matroid $M_{i}=(\mathcal{I},E_{i})$, and private values $v^{i}_{1},\ldots,v^{i}_{m}\in\mathbb{R}_{\geq 0}$ where agent’s $i$ valuation function $v_{i}:2^{\mathcal{I}}\rightarrow\mathbb{R}_{\geq 0}$ satisfies $v_{i}(S)=\max_{T\in E_{i}}\sum_{j\in S\cap T}v^{i}_{j}$. For this setting, we can use an adaptation of Mechanism 1 where instead of using a subroutine that gives the bundle with probability $\frac{q}{2^{\ell}}$, we can implement Step 1 by allocating each item independently with probability $\frac{1}{2^{\ell}}$ to one of its recipients of the copies. This modified mechanism is truthful in expectation since for weighted matroid functions (or even additive subject to downward-closed constraints) it holds that for every set $A_{i}^{\prime}$ allocated to agent $i$ at Step 1 of Mechanism 1 it holds that $g(A_{i}^{\prime})\in E_{i}$ and the agent’s valuation is additive on the potential sets allocated to him. For this to hold, the mechanism needs to know the matroids $M_{i}$ (if the matroids are private information, and the agents report them, then the proof of the truthfulness in expectation breaks).
Since weighted matroid functions are gross-substitutes, and since weighted matorid functions are closed [^7] By Claim 4, matroid rank functions are closed under the transformation that creates copies, and by similar arguments the same holds with respect to weighted matroid functions. under the operation of adding copies of items we get that we can extend the $O(\log(n))$ approximation with respect to the welfare also to this case. We leave as an open direction to extend our framework (or variants of it) beyond the settings studied in this paper, or in general devise truthful mechanisms with optimal surplus guarantees for other classes of functions (e.g., submodular, XOS, subadditive, supermodular, and additive subject to downward-closed constraints).
Appendix A Omitted Proofs
A.1 Proof of Claim 14
Proof of Claim 14.
$\sum_{i\in\mathcal{N}}{p_{i}(\mathcal{N},q,\mathcal{I})}\leq\sum_{i\in\mathcal{N}}{SW(\mathcal{N},q,\mathcal{I}+A_{i}(\mathcal{N},q,\mathcal{I}))-SW(\mathcal{N},q,\mathcal{I})}$ (8)
where the first inequality is due to the fact that the resulting allocation from giving each agent exactly half of what she receives in $SW(\mathcal{N},2q,\mathcal{I})$ is a feasible allocation when the item input is $\mathcal{I}/2$ and capacity is $q$, the first equality is by telescoping sum, the second inequality is due to the fact that by Claims 12 and 13 we have that the social welfare function is concave, and the final inequality is also due to this fact. Combining Equations (8) and (9) then completes the proof of the claim. ∎

### v3

@@ Appendix B Improved Approximations for Two Agents
In light of our results in the previous sections, which show that the surplus guarantees achievable in several well-studied multi-parameter auction settings with many agents are no worse than the surplus guarantees achievable in the canonical single-item setting with many agents, a natural question is whether there exist settings which are more challenging from an approximation standpoint than the single-item setting. To explore this question further, we now turn to the interesting special case of two agents. In previous work of [hartline2008optimal], it was demonstrated that there exists a mechanism for two agents and a single item that achieves surplus which is a $3/2$-approximation to the optimal social welfare [^8] More precisely, they demonstrate that their mechanism achieves a $3/2$-approximation to a weaker benchmark $\mathcal{G}$ which we discuss in Subsection B.1. The proof of this guarantee, however, implies a $3/2$-approximation to the stronger benchmark of optimal social welfare, which is of interest to the majority of our paper.. We generalize this result with Mechanism 3 which, as we demonstrate in Theorem 16, obtains a $3/2$-approximation guarantee in the case of two agents with arbitrary monotone valuation functions over an item set $\mathcal{I}$.
Input: An item set $\mathcal{I}$, a valuation profile of two agents $v_{1},v_{2}$
@@ 2 With probability $1/3$: Allocate the grand bundle of items to bidder $1$ and charge nothing
Consider two agents drawing their values i.i.d. from a standard exponential distribution. As demonstrated by [hartline2008optimal], the mechanism that maximizes the surplus for such a setting randomly allocates the item to one of the two agents and the expected surplus of such a mechanism is $1$. On the other hand, the expected social welfare is equal to the expectation of the maximum of two random variables $X_{1}$ and $X_{2}$ drawn independently from an exponential distribution. We may compute this as
@@
B.1 Alternative Benchmarks
Recall that the mechanisms in Sections 2 and 3 achieve surplus guarantees which are optimal approximations to the social welfare. However, these guarantees are asymptotic and logarithmic, raising the question of whether or not there exist tighter benchmarks that can better separate the performance of the two mechanisms. As all prior-free mechanisms are anonymous (i.e., cannot treat differently agents which are, a priori, identical), we would like to compare the performance of a prior-free mechanism to the surplus achieved by the best anonymous, truthful mechanism. This observation led to the development of the benchmark $\mathcal{G}(\vec{v})$ in [hartline2008optimal]. Formally, $\mathcal{G}(\vec{v})=\sup_{F}\text{Opt}_{F}(\vec{v})$ where $\text{Opt}_{F}(\vec{v})$ is the surplus of the Bayesian optimal mechanism when all bidders draw their value from a known distribution $F$ and the realization of values is $\vec{v}$. hartline2008optimal provide a prior-free mechanism which obtains a $O(1)$-approximation to $\mathcal{G}$ in the case of unit-demand bidders and identical items. In addition, they raise an open question of determining the best-possible approximation ratio when the benchmark is $\mathcal{G}$ “even in the two agent, one unit special case” [hartline2008optimal]. We resolve this open question below.
@@ 4     With the remaining probability allocate the item to bidder $2$ and charge bidder $2$ a price of 00
Mechanism 4 is truthful-in-expectation, ex-post individually rational, and obtains surplus utility that is a $5/4$-approximation to $\mathcal{G}(\vec{v})$.
@@
Recall that in single-parameter environments, a mechanism is truthful if it has a monotone allocation rule and charges payments corresponding to the identity given in [myerson1981optimal]. By observation, it is clear that Mechanism 4 is monotone, i.e., the probability a bidder is allocated a good is (weakly) increasing with her bid (holding the other bid constant). As such, it remains to verify that the prices charged by the mechanism are the threshold payments (i.e., those implied by the payment identity of myerson1981optimal). Consider varying the value $v_{i}$ of bidder $i$ when the value of the other bidder $j$ is fixed to $v_{j}$. When $0\leq v_{i}<\frac{1}{3}v_{j}$ the probability that bidder $i$ receives the good is $0.2$ and she pays 00, as expected by the payment identity. When $\frac{1}{3}v_{j}\leq v_{i}\leq 3v_{j}$ the probability that $i$ receives the good is $0.5$. As such, by the payment identity, $i$ should be charged $\frac{1}{3}v_{j}\cdot(0.5-0.2)=\frac{1}{10}v_{j}$ and this is precisely the price the mechanism charges (normalized by the probability to be allocated which is half). Finally, when $3v_{j}<v_{i}$ the probability that $i$ receives the good is $0.8$. But then, by the payment identity, $i$ should be charged $3v_{j}\cdot(0.8-0.5)+\frac{1}{10}v_{j}=v_{j}$ which is precisely what the mechanism charges (again, normalized by the allocation probability $\frac{4}{5}$).
Now, we turn to the analysis of the mechanism’s utility. Index the bidders in decreasing order of value such that we have $v_{1}\geq v_{2}$. As demonstrated in [hartline2008optimal], in the special case of two agents and a single item, $\mathcal{G}(\vec{v})$ is equal to the better of performing a Vickrey auction or random allocation. In other words, we have that $\mathcal{G}(\vec{v})=\max\left\{v_{1}-v_{2},\frac{v_{1}+v_{2}}{2}\right\}$ [^9] We note that in the proof of Proposition 4.1 in [hartline2008optimal], after arguing that $\mathcal{G}(\vec{v})$ is the better of a Vickrey auction or random allocation in the case of two agents, they incorrectly compute that $\mathcal{G}(\vec{v})=\max\left\{v_{1}-\frac{v_{2}}{2},\frac{v_{1}+v_{2}}{2}\right\}$. This leads them to give an incorrect lower bound of $4/3>5/4$ on the approximation achievable by any mechanism.
@@
Our proof proceeds very similarly to the proof of Proposition 4.1 in [hartline2008optimal] (essentially, correcting an arithmetic error that propagates therein). Consider two agents drawing their values i.i.d. from a standard exponential distribution. Again, the mechanism which maximizes the expected surplus for such a setting randomly allocates the item to one of the two agents and the expected surplus of such a mechanism is $1$. On the other hand, the expected value of $\mathcal{G}(\vec{v})$ is